# Hopf orbits and the first ECH capacity

### v2

@@ 3.1. Embedded contact homology
To begin, let $Y$ be a closed three-manifold and let $\lambda$ be a nondegenerate contact form on $Y$. We now review the definition of embedded contact homology ${\rm ECH}(Y,\lambda)$. For simplicity we further assume that $Y$ is a homology sphere, which is sufficient for our purposes; for more details and the case of general three-manifolds see e.g. [ind_rev, 26].
@@
If $\alpha=\alpha_{1}^{m_{1}}\cdots\alpha_{k}^{m_{k}}$ is an ECH generator, its ECH index is the integer $I(\alpha)$ defined by [^6] To compare with the notation in other papers: $I(\alpha)$ here corresponds to $I(\alpha,\emptyset)$ in [ind_rev, Def. 2.15], and to $I(\alpha,\emptyset,Z)$ in [26, Eq. (3.4)], where $Z$ is the unique relative homology class of surface in $Y$ with boundary $\alpha$.
(3.1) $I(\alpha)=\sum_{i}\left(m_{i}c_{\tau}(\alpha_{i})+m_{i}^{2}Q_{\tau}(\alpha_{i})\right)+\sum_{i\neq j}m_{i}m_{j}\operatorname{link}(\alpha_{i},\alpha_{j})+\sum_{i}\sum_{l=1}^{m_{i}}{\rm CZ}_{\tau}(\alpha_{i}^{l}).$
Here $\tau$ is any trivialization of $\xi$ over the Reeb orbits $\alpha_{i}$; the ECH index $I(\alpha)$ does not depend on this choice, although the terms on the right hand side of (3.1) do. If $\gamma$ is a simple Reeb orbit, then $c_{\tau}(\gamma)$ denotes the relative first Chern class of $\xi$ with respect to $\tau$ over a surface bounded by $\gamma$, and $Q_{\tau}(\gamma)$ denotes the relative self-intersection number with respect to $\tau$ of a surface bounded by $\gamma$; see [ind_rev, 26] for details. Also, $\operatorname{link}$ denotes the linking number, and $\gamma^{l}$ denotes the Reeb orbit that covers $\gamma$ with multiplicity $l$.
@@
(3.3) $I(\alpha)=\sum_{i}m_{i}^{2}{\rm sl}(\alpha_{i})+\sum_{i\neq j}m_{i}m_{j}\operatorname{link}(\alpha_{i},\alpha_{j})+\sum_{i}\sum_{l=1}^{m_{k}}{\rm CZ}(\alpha_{i}^{l}).$
@@
Next, choose a generic $d\lambda$-compatible almost complex structure on $\xi$, and let $\widetilde{J}$ denote the corresponding almost complex structure on ×$\mathbb{R}\times Y$ as in §2.3. If $\alpha$ and $\beta$ are ECH generators, define $\mathcal{M}^{J}(\alpha,\beta)$ to be the moduli space of $\widetilde{J}$-holomorphic currents [^8] A “holomorphic current” is a finite linear combination of somewhere injective holomorphic curves with positive integer multiplicities; see [26, §3.1] for details. in ×$\mathbb{R}\times Y$ asymptotic to $\alpha$ as the $\mathbb{R}$ coordinate goes to +$\infty$+ ∞ and asymptotic to $\beta$ as the $\mathbb{R}$ coordinate goes to $\infty$- ∞. If $J$ is generic, then when $I(\alpha)-I(\beta)=1$, the moduli space $\mathcal{M}^{J}(\alpha,\beta)/\mathbb{R}$ is finite, where $\mathbb{R}$ acts by translation of the $\mathbb{R}$ coordinate on ×$\mathbb{R}\times Y$; see [26, §5.3].
@@
$\partial_{J}:ECC_{*}(Y,\lambda)\longrightarrow ECC_{*-1}(Y,\lambda)$
@@
where $\beta$ is an ECH generator and ##\## denotes the mod 2 count.
It is shown in [HT1, Thm. 7.20] that $\partial_{J}^{2}=0$. We denote the homology of the chain complex $(ECC_{*}(Y,\lambda),\partial_{J})$ by $ECH_{*}(Y,\lambda,J)$. It was shown by Taubes [taubes_iso] that $ECH_{*}(Y,\lambda,J)$ is canonically isomorphic to a version of Seiberg-Witten Floer cohomology as defined by Kronheimer-Mrowka [32]. As explained in [31, Thm. 1.3], this implies that $ECH_{*}(Y,\lambda,J)$ in fact depends only on the pair $(Y,\xi)$, so we can denote it by $ECH_{*}(Y,\xi)$.
@@
See e.g. the calculation in [26, §3.7].
There is some additional structure on ECH which we will need. First of all, the empty set is a valid ECH generator, with grading $I(\emptyset)=0$. For any $J$ we have $\partial_{J}\emptyset=0$, because the ECH differential decreases action, as explained in §3.2 below. It follows from the properties of ECH cobordism maps in [31, Thm. 1.9] that the homology class $[\emptyset]\in\operatorname{ECH}_{0}(Y,\lambda,J)$ gives a well-defined invariant
@@
which counts $\widetilde{J}$-holomorphic currrents passing through $(0,y)$ where $y\in Y$ is a point not on any Reeb orbit. See [26, §3.8] for details about why the $U$ map is well defined. In the example (3.4), the $U$ map on $\operatorname{ECH}_{*}(S^{3},\xi_{0})$ is an isomorphism whenever *0$\neq$ 0* ≠ 0.
@@ 3.2. ECH capacities
It is shown in [31, Thm. 1.3] that the filtered ECH depends only on $Y,\lambda,L$ and not on the choice of $J$. In addition, inclusion of chain complexes inducs a well-defined chain map
@@
We first prove that $c_{\rm Hopf}(X)\geq c_{1}^{{\rm ECH}}(X)$. One can quickly deduce this step from the result of Edtmair [gss_edtmair] that $c_{Z}(X)\leq\mathcal{A}_{\operatorname{Hopf}}(X)$, but we will give an independent proof.
By Theorem 1.12, there exists a Hopf orbit $\gamma\in\mathcal{P}_{\operatorname{Hopf}}(\partial X)$ satisfying $c_{\operatorname{Hopf}}(X)=\mathcal{A}(\gamma)$ and ${\rm CZ}(\gamma)=3$. It follows from equation (3.3) that the ECH index $I(\gamma)=2$.
Now choose a generic almost complex structure $J$ as needed to define the ECH differential $\partial_{J}$, and choose a point $y\in Y$ as needed to define the chain map $U_{J,y}$. We claim that $\partial_{J}\gamma=0$ and $U_{J,y}\gamma=\emptyset$. Assuming this claim, it follows from the nontriviality of the class []delimited-[][$\emptyset$][ ∅ ] that $\gamma$ is a cycle in the ECH chain complex representing the generator of $\operatorname{ECH}_{2}(Y,\xi_{0})$, so it follows from Definition 3.8 that $c_{1}^{\operatorname{ECH}}(X)\leq\mathcal{A}(\gamma)$, completing the proof of Step 1.
To prepare to prove the claim, recall from Remark 1.9 that since $X$ is dynamically convex and $\gamma$ is a Hopf orbit, it follows that $\gamma$ is the boundary of a disk-like global surface of section $D\subset Y$. In particular, every Reeb orbit other than $\gamma$ has positive linking number with $\gamma$. Moreover, the construction of the global surface of section in [hryn_jsg] obtains $\operatorname{int}(D)$ as the projection to $Y$ of a $J$-holomorphic plane in ×$\mathbb{R}\times Y$ with a positive puncture at $\gamma$.
@@
It follows from the asymptotics in Theorem 2.4 that if $a_{0}$ is sufficiently large, then $C$ is transverse to {±$\{\pm a_{0}\}\times Y$, and
@@
(3.6) $\mathrm{link}(\eta_{+},\gamma)-\mathrm{link}(\eta_{-},\gamma)=\#(C\cap([-a_{0},a_{0}]\times\gamma))\geq 0,$
where ##\## denotes the algebraic count of intersections, which is nonnegative by intersection positivity of holomorphic curves.
@@
Combining the inequalities (3.6), (3.7), and (3.8) gives $-1\geq 0$, a contradiction.
@@
Combining (3.9), (3.10), and (3.11), we obtain #$\#(C\cap C^{\prime})\leq 0$. By intersection positivity, it follows that #$\#(C\cap C^{\prime})=0$ and $C\cap C^{\prime}=\emptyset$ as desired.
@@
We now show that any cycle in the chain complex $(ECC_{*}(Y,\lambda),\partial_{J})$ that is homologous to $\emptyset$ must include $\emptyset$ as a summand. For this purpose it suffices to show that if $\alpha$ is an ECH generator with $I(\alpha)=1$, then $\mathcal{M}^{J}(\alpha,\emptyset)=\emptyset$. One can prove this using the Fredholm index, but to prepare for the following step we will give a different proof using the “$J_{0}$ index” defined in [ind_rev, §6.1].
@@
(3.12) $J_{0}(\alpha)=I(\alpha)-\sum_{i=1}^{k}{\rm CZ}(\alpha_{i}^{m_{i}}).$
@@
with equality only if $\mathcal{C}$ consists of a holomorphic plane, possibly together with some trivial cylinders. This is a special case of [26, Prop. 5.8].
@@
On the other hand $J_{0}(\emptyset)=0$. Then if $\mathcal{M}^{J}(\alpha,\emptyset)\neq\emptyset$, the inequality (3.13) with $\beta=\emptyset$ gives $-2\geq-1$, a contradiction.
@@
We now prove that $c_{\rm Hopf}(X)\leq c_{1}^{{\rm ECH}}(X)$.
@@
By the inequality (3.13), we have $J_{0}(x_{j})\geq-1$. By equation (3.12) and dynamical convexity, $x_{j}$ consists of a single simple Reeb orbit $\gamma$ with multiplicity $1$, and $\operatorname{CZ}(\gamma)=3$. Since $I(\gamma)=2$, it follows from equation (3.3) that $\operatorname{sl}(\gamma)=-1$. Since equality holds in (3.13), $\mathcal{C}$ is a holomorphic plane positively asymptotic to $\gamma$. By [30, Prop. 3.4], the projection of $\mathcal{C}$ to $Y$ is an embedding [^9] The statement of [30, Prop. 3.4] requires that $\gamma$ is elliptic, but the proof works just as well if $\gamma$ is negative hyperbolic, so it applies given that $\operatorname{CZ}(\gamma)=3$. The basic ideas in the proof go back to [props2], and a more general statement is proved in [sief_int, Theorem 2.6]. One can further show that the projection of $\mathcal{C}$ to $Y$ is the interior of a global surface of section for the Reeb flow, but we do not need this.. Thus $\gamma$ is unknotted in $Y$, so $\gamma$ is a Hopf orbit. Then $c_{\operatorname{Hopf}}(X)\leq\mathcal{A}(\gamma)$, and we know from the previous paragraph that $\mathcal{A}(\gamma)\leq c_{1}^{\operatorname{ECH}}(X)$.

### v3

@@ 3.1. Embedded contact homology
To begin, let $Y$ be a closed three-manifold and let $\lambda$ be a nondegenerate contact form on $Y$. We now review the definition of embedded contact homology $\operatorname{ECH}(Y,\lambda)$. For simplicity we further assume that $Y$ is a homology sphere, which is sufficient for our purposes; for more details and the case of general three-manifolds see e.g. [ind_rev, 32].
@@
If $\alpha=\alpha_{1}^{m_{1}}\cdots\alpha_{k}^{m_{k}}$ is an ECH generator, its ECH index is the integer $I(\alpha)$ defined by [^5] To compare with the notation in other papers: $I(\alpha)$ here corresponds to $I(\alpha,\emptyset)$ in [ind_rev, Def. 2.15], and to $I(\alpha,\emptyset,Z)$ in [32, Eq. (3.4)], where $Z$ is the unique relative homology class of surface in $Y$ with boundary $\alpha$.
(3.1) $I(\alpha)=\sum_{i}\left(m_{i}c_{\tau}(\alpha_{i})+m_{i}^{2}Q_{\tau}(\alpha_{i})\right)+\sum_{i\neq j}m_{i}m_{j}\operatorname{link}(\alpha_{i},\alpha_{j})+\sum_{i}\sum_{l=1}^{m_{i}}\operatorname{CZ}_{\tau}(\alpha_{i}^{l}).$
Here $\tau$ is any trivialization of $\xi$ over the Reeb orbits $\alpha_{i}$; the ECH index $I(\alpha)$ does not depend on this choice, although the terms on the right hand side of (3.1) do. If $\gamma$ is a simple Reeb orbit, then $c_{\tau}(\gamma)$ denotes the relative first Chern class of $\xi$ with respect to $\tau$ over a surface bounded by $\gamma$, and $Q_{\tau}(\gamma)$ denotes the relative self-intersection number with respect to $\tau$ of a surface bounded by $\gamma$; see [ind_rev, 32] for details. Also, $\operatorname{link}$ denotes the linking number, and $\gamma^{l}$ denotes the Reeb orbit that covers $\gamma$ with multiplicity $l$.
@@
(3.3) $I(\alpha)=\sum_{i}m_{i}^{2}\operatorname{sl}(\alpha_{i})+\sum_{i\neq j}m_{i}m_{j}\operatorname{link}(\alpha_{i},\alpha_{j})+\sum_{i}\sum_{l=1}^{m_{k}}\operatorname{CZ}(\alpha_{i}^{l}).$
Here $\operatorname{CZ}$ stands for the Conley-Zehnder index taken with respect to $\tau$.
@@
Next, choose a generic $d\lambda$-compatible almost complex structure on $\xi$, and let $\widetilde{J}$ denote the corresponding almost complex structure on $\mathbb{R}\times Y$ as in §2.3. If $\alpha$ and $\beta$ are ECH generators, define $\mathcal{M}^{J}(\alpha,\beta)$ to be the moduli space of $\widetilde{J}-$holomorphic currents [^7] A “holomorphic current” is a finite linear combination of somewhere injective holomorphic curves with positive integer multiplicities; see [32, §3.1] for details. in $\mathbb{R}\times Y$ asymptotic to $\alpha$ as the $\mathbb{R}$ coordinate goes to $\infty$ and asymptotic to $\beta$ as the $\mathbb{R}$ coordinate goes to -$\infty$. If $J$ is generic, then when $I(\alpha)-I(\beta)=1$, the moduli space $\mathcal{M}^{J}(\alpha,\beta)/\mathbb{R}$ is finite, where $\mathbb{R}$ acts by translation of the $\mathbb{R}$ coordinate on $\mathbb{R}\times Y$; see [32, §5.3].
@@
$\partial_{J}:\operatorname{ECC}_{*}(Y,\lambda)\longrightarrow\operatorname{ECC}_{*-1}(Y,\lambda)$
@@
where $\beta$ is an ECH generator and #\# denotes the mod 2 count.
It is shown in [HT1, Thm. 7.20] that $\partial_{J}^{2}=0$. We denote the homology of the chain complex $(\operatorname{ECC}_{*}(Y,\lambda),\partial_{J})$ by $\operatorname{ECH}_{*}(Y,\lambda,J)$. It was shown by Taubes [taubes_iso] that $\operatorname{ECH}_{*}(Y,\lambda,J)$ is canonically isomorphic to a version of Seiberg-Witten Floer cohomology as defined by Kronheimer-Mrowka [38]. As explained in [37, Thm. 1.3], this implies that $\operatorname{ECH}_{*}(Y,\lambda,J)$ in fact depends only on the pair $(Y,\xi)$, so we can denote it by $\operatorname{ECH}_{*}(Y,\xi)$.
@@
See e.g. the calculation in [32, §3.7].
There is some additional structure on ECH which we will need. First of all, the empty set is a valid ECH generator, with grading $I(\emptyset)=0$. For any $J$ we have $\partial_{J}\emptyset=0$, because the ECH differential decreases action, as explained in §3.2 below. It follows from the properties of ECH cobordism maps in [37, Thm. 1.9] that the homology class $\emptyset]\in\operatorname{ECH}_{0}(Y,\lambda,J)$ gives a well-defined invariant
@@
which counts $\widetilde{J}-$holomorphic currrents passing through $(0,y)$ where $y\in Y$ is a point not on any Reeb orbit. See [32, §3.8] for details about why the $U$ map is well defined. In the example (3.4), the $U$ map on $\operatorname{ECH}_{*}(S^{3},\xi_{0})$ is an isomorphism whenever $\neq 0$.
@@ 3.2. ECH capacities
It is shown in [37, Thm. 1.3] that the filtered ECH depends only on $Y,\lambda,L$ and not on the choice of $J$. In addition, inclusion of chain complexes inducs a well-defined chain map
@@
We first prove that $c_{\rm Hopf}(X)\geq c_{1}^{\operatorname{ECH}}(X)$. One can quickly deduce this step from the result of Edtmair [13] that $c_{Z}(X)\leq\mathcal{A}_{\operatorname{Hopf}}(X)$, but we will give an independent proof.
By Theorem 1.9, there exists a Hopf orbit $\gamma\in\mathcal{P}_{\operatorname{Hopf}}(\partial X)$ satisfying $c_{\operatorname{Hopf}}(X)=\mathcal{A}(\gamma)$ and $\operatorname{CZ}(\gamma)=3$. It follows from equation (3.3) that the ECH index $I(\gamma)=2$.
Now choose a generic almost complex structure $J$ as needed to define the ECH differential $\partial_{J}$, and choose a point $y\in Y$ as needed to define the chain map $U_{J,y}$. We claim that $\partial_{J}\gamma=0$ and $U_{J,y}\gamma=\emptyset$. Assuming this claim, it follows from the nontriviality of the class $\emptyset$] that $\gamma$ is a cycle in the ECH chain complex representing the generator of $\operatorname{ECH}_{2}(Y,\xi_{0})$, so it follows from Definition 3.8 that $c_{1}^{\operatorname{ECH}}(X)\leq\mathcal{A}(\gamma)$, completing the proof of Step 1.
To prepare to prove the claim, recall from Remark 1.6 that since $X$ is dynamically convex and $\gamma$ is a Hopf orbit, it follows that $\gamma$ is the boundary of a disk-like global surface of section $D\subset Y$. In particular, every Reeb orbit other than $\gamma$ has positive linking number with $\gamma$. Moreover, the construction of the global surface of section in [hryn_jsg] obtains $\operatorname{int}(D)$ as the projection to $Y$ of a $J$-holomorphic plane in $\mathbb{R}\times Y$ with a positive puncture at $\gamma$.
@@
It follows from the asymptotics in Theorem 2.5 that if $a_{0}$ is sufficiently large, then $C$ is transverse to $\{\pm a_{0}\}\times Y$, and
@@
(3.6) $\operatorname{link}(\eta_{+},\gamma)-\operatorname{link}(\eta_{-},\gamma)=\#(C\cap([-a_{0},a_{0}]\times\gamma))\geq 0,$
where #\# denotes the algebraic count of intersections, which is nonnegative by intersection positivity of holomorphic curves.
@@
Combining the inequalities (3.6), (3.7), and (3.8) gives $1\geq 0$, a contradiction.
@@
Combining (3.9), (3.10), and (3.11), we obtain $(C\cap C^{\prime})\leq 0$. By intersection positivity, it follows that $(C\cap C^{\prime})=0$ and $C\cap C^{\prime}=\emptyset$ as desired.
@@
We now show that any cycle in the chain complex $(ECC_{*}(Y,\lambda),\partial_{J})$ that is homologous to $\emptyset$ must include $\emptyset$ as a summand. For this purpose it suffices to show that if $\alpha$ is an ECH generator with $I(\alpha)=1$, then $\mathcal{M}^{J}(\alpha,\emptyset)=\emptyset$. One can prove this using the Fredholm index, but to prepare for the following step we will give a different proof using the $J_{0}$ index” defined in [ind_rev, §6.1].
@@
(3.12) $J_{0}(\alpha)=I(\alpha)-\sum_{i=1}^{k}\operatorname{CZ}(\alpha_{i}^{m_{i}}).$
@@
with equality only if $\mathcal{C}$ consists of a holomorphic plane, possibly together with some trivial cylinders. This is a special case of [32, Prop. 5.8].
@@
On the other hand $J_{0}(\emptyset)=0$. Then if $\mathcal{M}^{J}(\alpha,\emptyset)\neq\emptyset$, the inequality (3.13) with $\beta=\emptyset$ gives $2\geq-1$, a contradiction.
@@
We now prove that $c_{\rm Hopf}(X)\leq c_{1}^{\operatorname{ECH}}(X)$.
@@
By the inequality (3.13), we have $J_{0}(x_{j})\geq-1$. By equation (3.12) and dynamical convexity, $x_{j}$ consists of a single simple Reeb orbit $\gamma$ with multiplicity $1$, and $\operatorname{CZ}(\gamma)=3$. Since $I(\gamma)=2$, it follows from equation (3.3) that $\operatorname{sl}(\gamma)=-1$. Since equality holds in (3.13), $\mathcal{C}$ is a holomorphic plane positively asymptotic to $\gamma$. By [36, Prop. 3.4], the projection of $\mathcal{C}$ to $Y$ is an embedding [^8] The statement of [36, Prop. 3.4] requires that $\gamma$ is elliptic, but the proof works just as well if $\gamma$ is negative hyperbolic, so it applies given that $\operatorname{CZ}(\gamma)=3$. The basic ideas in the proof go back to [props2], and a more general statement is proved in [sief_int, Theorem 2.6]. One can further show that the projection of $\mathcal{C}$ to $Y$ is the interior of a global surface of section for the Reeb flow, but we do not need this.. Thus $\gamma$ is unknotted in $Y$, so $\gamma$ is a Hopf orbit. Then $c_{\operatorname{Hopf}}(X)\leq\mathcal{A}(\gamma)$, and we know from the previous paragraph that $\mathcal{A}(\gamma)\leq c_{1}^{\operatorname{ECH}}(X)$.